OpenStudy (anonymous):

Ok, time for me to quit for the night, but I'll leave y'all with this to scratch your heads about: Infinite sum of 1+2+4+8+16+32+ . . . = -1. Proof: Multiply series by 1 (which should not change the sum); however, I'm going to express 1 as (2-1). 2-1=1. (2-1)(1+2+4+8+16+...) =(2+4+8+16+32+...)+(-1-2-4-8-16-...) All the opposites cancel out and all that is left is -1. Pretty screwy, huh?

OpenStudy (anonymous):

doesnt work like this answer is still infinity

OpenStudy (anonymous):

in a nutshell you subtracting an n+1 term summation from an n term summation since both summations are to infinity so they must have the same amount of terms thus n+1=n 1=0

OpenStudy (anonymous):

ive asked my professor this before sorry i dont remember the full explanation

OpenStudy (anonymous):

The explanation I give most people is "infinity is weird." ;-)

OpenStudy (anonymous):

Of course it's a bogus proof, but it's fun to show people to give them that paradox feeling.

OpenStudy (anonymous):

i will represent the infinite summation as x x= infinity 2x and 1x 2x-1x=x(2-1)=x(1) its not nice to trick people

OpenStudy (anonymous):

LOL, what do you mean "it's not nice to trick people." It's a puzzle, the object is to find out what's wrong with it. You obviously weren't tricked.

OpenStudy (anonymous):

misleadddd*

OpenStudy (anonymous):

Please. Would you have felt better if I prefaced it with the statement, "The following is a bogus proof, try to find what's wrong with it."? I'm sure it's pretty clear to anyone that it's not serious.

OpenStudy (anonymous):

Buncha grouches 'round here . . .